OpenStudy (anonymous):

y-2x+3=0 in funcyion notation

OpenStudy (anonymous):

can any body find this solution

OpenStudy (anonymous):

y-2x+3=0 in funcyion notation

OpenStudy (anonymous):

hi

OpenStudy (javk):

ok so...function notation means that you need to rearrange your eqn so that you have y on one side and all the x's on the other y=2x-3 remember that y=f(x) so... f(x)=2x-3 and that should be your ans